# Towards Open Respiratory Acoustic Foundation Models: Pretraining and Benchmarking

### v1

@@ Abstract
Respiratory audio, such as coughing and breathing sounds, has predictive power for a wide range of healthcare applications, yet is currently under-explored. The main problem for those applications arises from the difficulty in collecting large labeled task-specific data for model development. Generalizable respiratory acoustic foundation models pretrained with unlabeled data would offer appealing advantages and possibly unlock this impasse. However, given the safety-critical nature of healthcare applications, it is pivotal to also ensure openness and replicability for any proposed foundation model solution. To this end, we introduce OPERA, an OPEn Respiratory Acoustic foundation model pretraining and benchmarking system, as the first approach answering this need. We curate large-scale respiratory audio datasets ($\sim$136K samples, 440 hours), pretrain three pioneering foundation models, and build a benchmark consisting of 19 downstream respiratory health tasks for evaluation. Our pretrained models demonstrate superior performance (against existing acoustic models pretrained with general audio on 16 out of 19 tasks) and generalizability (to unseen datasets and new respiratory audio modalities). This highlights the great promise of respiratory acoustic foundation models and encourages more studies using OPERA as an open resource to accelerate research on respiratory audio for health. The system is accessible from https://github.com/evelyn0414/OPERA.
@@ 1 Introduction
There is no open respiratory acoustic foundation model, impeding the field’s growth and understanding. Existing open-source acoustic models like AudioMAE [36] and CLAP [18] are pretrained on general audio event datasets such as YouTube audio, containing very few (around 0.3%) respiratory sounds [39, 25]. These models may not be able to effectively capture the subtle nuances of respiratory sounds, which can vary in abrupt bursts, aperiodic components, and frequency distributions, particularly across different health conditions [50]. Although a model pretrained on respiratory sounds has been recently presented [7], it is not open-source, making it hard to analyze, replicate, or compare its workings. The insights on how to effectively train respiratory acoustic foundation models also remain limited.
@@
Figure 1: System overview of OPERA. After data curation, respiratory acoustic foundation models (Encoder) are pretrained and then evaluated on various downstream health tasks.
To mitigate these gaps, in this paper, we put forward OPERA, an OPEn Respiratory Acoustic foundation model pretraining and benchmarking system (Figure 1). It curates unlabeled respiratory audio datasets, pretrain three pioneering foundation models, and evaluates them against existing pretrained acoustic models across various applications. Specifically, our contributions are:
  • We curate a unique large-scale ($\sim$136K samples, 440 hours), multi-source (5 datasets), multi-modal (breathing, coughing, and lung sounds) and publicly available (or available on request) respiratory audio dataset for foundation model pretraining, orders of magnitude larger than the number of respiratory audio samples in datasets used for training existing open acoustic models.
  • We pretrain 3 foundation models with the curated unlabeled data using the most common self-supervised approaches (a contrastive learning-based transformer, a contrastive learning-based CNN model, and a generatively pretrained transformer) to study the effect of the training designs.
@@
  • We benchmark the performance of our 3 foundation models, one commonly used acoustic feature set, and 3 open pretrained acoustic models on these tasks as a starting point for future exploration.
Extensive experiments demonstrate that our foundation models outperform the models pretrained with general audio on 16 out of 19 benchmark tasks, confirming the power and promise of dedicated respiratory acoustic foundation models. Results also show that our models are generalizable across multiple downstream tasks, including new datasets and unseen respiratory audio modalities. This is a critical advancement towards realizing the potential of respiratory sounds as a mainstream technique for health monitoring.
@@
In summary, this paper introduces the first open-source respiratory acoustic foundation model pretraining and benchmarking system. This represents a critical first step towards comprehensive and reproducible audio foundation models for health: future foundation model research can leverage our system as an experimental resource, and application studies can take advantage of our foundation models as feature extractors. This can facilitate progress in both machine learning and healthcare. These efforts will extend current machine learning capabilities, now able to see (via vision) and read (via natural languages), to also listen to (via audio) our health.
@@ 2.1 Pretraining in Acoustic Modeling
It is also worth noting that these open models are pretrained on general audio event datasets such as AudioSet [25], FSD50K [22], and FreeSound [23], which contains few samples of respiratory-related audio. For instance, AudioSet’s 2 million clips include only 2334 snoring, 871 cough, 834 breathing, and 1200 sneeze clips, making up only 0.3% of the total. In face of this issue, we curate large-scale respiratory audio datasets to pretrain our foundation models for comparison.
@@
A recent work, HeAR [7], curated millions of respiratory audio clips from YouTube videos to pretrain a foundation model using a generative SSL approach. However, neither the data nor the model are publicly available, resulting in a lack of transparency and reproducibility. Limited exploration has been conducted on the reasoning behind the chosen SSL method for various downstream tasks. Our work investigates, for the first time, open pretraining respiratory acoustic foundation models to provide a better understanding of their limits and their potential.
@@ 3 System Overview
As shown in Figure 1, OPERA comprises three main components: data curation (including unlabeled data for pretraining and labeled data for evaluation), general-purpose pretraining to develop acoustic foundation models (Encoder), and a benchmark comparing the pretrained models on various downstream tasks.
@@ 5.1 Benchmark Datasets and Tasks Setup
Dataset ID Task Modality #Sam. (#Sub.) Data Distribution UK COVID-19 [13] T1 Covid / Non-covid Exhalation 2500 (2500) 840 / 1660 T2 Covid / Non-covid Cough 2500 (2500) 840 / 1660 COVID-19 Sounds [70] T3 Symptomatic / Healthy Breath 4138 (3294) 2029 / 2109 T4 Symptomatic / Healthy Cough 4138 (3294) 2029 / 2109 CoughVID [48] T5 Covid / Non-covid Cough 6175 (n/a) 547 / 5628 T6 Female / Male Cough 7263 (n/a) 2468 / 4795 ICBHI [52] T7 COPD / Healthy Lung sounds 828 (90) 793 / 35 Coswara [8] T8 Smoker / Non-smoker Cough 948 (n/a) 201 / 747 T9 Female / Male Cough 2496 (n/a) 759 / 1737 KAUH [24] T10 Obstructive / Healthy Lung sounds 234 (79) 129 / 105 Respiratory@TR [3] T11 COPD severity Lung sounds 504 (42) 72 / 60 / 84 / 84 / 204 SSBPR [71] T12 Body position recognition Snoring 7468 (20) 1638 / 1454 / 1269 / 1668 / 1439 MMlung [45] T13 FVC Deep breath 40 (40) 3.402 ± 1.032 L T14 FEV1 Deep breath 40 (40) 2.657 ± 0.976 L T15 FEV1/FVC Deep breath 40 (40) 0.808 ± 0.190 L T16 FVC O Vowels 40 (40) 3.402 ± 1.032 L T17 FEV1 O Vowels 40 (40) 2.657 ± 0.976 L T18 FEV1/FVC O Vowels 40 (40) 0.808 ± 0.190 L NoseMic [10] T19 Respiratory rate Breath 1297 (16) 13.915 ± 3.386 bpm
@@
Baselines. In addition to our pretrained models, we also include a commonly used acoustic feature set and three open pretrained acoustic models in this benchmark. They are Opensmile  [19] (Emobase acoutic feature set), VGGish [31] (supervised pretrained), AudioMAE [36] (self-supervised pretrained) and CLAP [18] (language-supervised pretrained). We consider these four methods as baselines to be distinguished from our pretrained models.
@@
Table 3: Mean reciprocal ranks on task groups (higher is better). The best model within each group is highlighted in pink and the second-best is highlighted in blue.
@@ 5.2 Experimental Results
RQ1. Can pretraining a foundation model with diverse unlabeled respiratory audio data lead to better performance than baselines designed for general audio?
@@
ID Task Abbr. Opensmile VGGish AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT T1 Covid (Exhale) 0.550 ± 0.015 0.580 ± 0.001 0.549 ± 0.001 0.565 ± 0.001 0.586 ± 0.008 0.551 ± 0.010 0.605 ± 0.001 ✓* T2 Covid (Cough) 0.649 ± 0.006 0.557 ± 0.005 0.616 ± 0.001 0.648 ± 0.003 0.701 ± 0.002 0.629 ± 0.006 0.677 ± 0.001 ✓* T3 Symptom (Breath) 0.571 ± 0.006 0.571 ± 0.003 0.583 ± 0.003 0.611 ± 0.006 0.603 ± 0.005 0.610 ± 0.004 0.613 ± 0.002 ✓* T4 Symptom (Cough) 0.633 ± 0.012 0.605 ± 0.004 0.659 ± 0.001 0.669 ± 0.002 0.680 ± 0.006 0.665 ± 0.001 0.673 ± 0.001 ✓* T5 Covid (Cough) 0.537 ± 0.011 0.538 ± 0.028 0.554 ± 0.004 0.599 ± 0.007 0.578 ± 0.001 0.566 ± 0.008 0.552 ± 0.003 T6 Gender (Cough) 0.677 ± 0.005 0.600 ± 0.001 0.628 ± 0.001 0.665 ± 0.001 0.795 ± 0.001 0.721 ± 0.001 0.735 ± $0$* T7 COPD (Lung) 0.579 ± 0.043 0.605 ± 0.077 0.886 ± 0.017 0.933 ± 0.005 0.855 ± 0.012 0.872 ± 0.011 0.741 ± 0.011 T8 Smoker (Cough) 0.534 ± 0.060 0.507 ± 0.027 0.549 ± 0.022 0.680 ± 0.009 0.685 ± 0.012 0.674 ± 0.013 0.650 ± 0.005 ✓* T9 Gender (Cough) 0.753 ± 0.008 0.606 ± 0.003 0.724 ± 0.001 0.742 ± 0.001 0.874 ± 0.000 0.801 ± 0.002 0.825 ± 0.001 ✓* T10 Obstructive (Lung) 0.636 ± 0.082 0.605 ± 0.036 0.616 ± 0.041 0.697 ± 0.004 0.722 ± 0.016 0.741 ± 0.014 0.703 ± 0.016 ✓* T11 COPD severity (Lung) 0.494 ± 0.054 0.590 ± 0.034 0.510 ± 0.021 0.636 ± 0.045 0.625 ± 0.038 0.683 ± 0.007 0.606 ± 0.015 ✓* T12 Position (Snoring) 0.772 ± 0.005 0.657 ± 0.002 0.649 ± 0.001 0.702 ± 0.001 0.781 ± 0.000 0.769 ± 0.000 0.742 ± 0.001 ✓*
@@
ID Task Abbr. Opensmile VGGish AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT T13 FVC (Breath) 0.985 ± 0.743 0.904 ± 0.568 0.900 ± 0.551 0.896 ± 0.542 0.924 ± 0.583 0.848 ± 0.607 0.892 ± 0.618 ✓* T14 FEV1 (Breath) 0.756 ± 0.721 0.839 ± 0.563 0.821 ± 0.590 0.840 ± 0.547 0.837 ± 0.563 0.834 ± 0.581 0.825 ± 0.560 T15 FEV1/FVC (Breath) 0.141 ± 0.185 0.131 ± 0.146 0.129 ± 0.146 0.134 ± 0.146 0.128 ± 0.140 0.132 ± 0.141 0.128 ± 0.141 ✓* T16 FVC (Vowel) 0.850 ± 0.592 0.895 ± 0.559 0.833 ± 0.588 0.883 ± 0.560 0.885 ± 0.553 0.761 ± 0.544 0.878 ± 0.550 ✓* T17 FEV1 (Vowel) 0.730 ± 0.497 0.842 ± 0.559 0.876 ± 0.561 0.859 ± 0.541 0.780 ± 0.542 0.830 ± 0.561 0.774 ± 0.554    * T18 FEV1/FVC (Vowel) 0.138 ± 0.166 0.130 ± 0.145 0.131 ± 0.141 0.137 ± 0.147 0.132 ± 0.140 0.136 ± 0.150 0.130 ± 0.138 ✓* T19 Breathing Rate 2.714 ± 0.902 2.605 ± 0.759 2.641 ± 0.813 2.650 ± 0.947 2.636 ± 0.858 2.525 ± 0.782 2.416 ± 0.885 ✓*
RQ2. Are the pretrained respiratory acoustic foundation models generalizable to new data?
It is crucial that foundation models can generalize to new and unseen data once developed. In our benchmark, we have 12 tasks formulated from unseen datasets (Task 8-19) and unseen respiratory audio modalities (Task 12, 16-18) not used for pretraining. Notably, our respiratory acoustic foundation models demonstrate good generalization capabilities, achieving the best performance on 5 out of 5 classification tasks and 5 out of 7 regression tasks. They are able to outperform the acoustic feature set and general audio pretrained models which are supposed to exhibit generalizability. Specifically, in Table 4, Task 8-12 all have an AUROC higher than 0.68. Comparing Task 6 and Task 9 with the same prediction target, the performance on unseen data (Task 9) is comparable. Therefore, our foundation models are generalizable, likely due to the minimal assumptions made during SSL pretraining.
RQ3. How to design SSL methods and model architectures of respiratory acoustic foundation models with different applications in mind?

### v2

@@ Abstract
Respiratory audio, such as coughing and breathing sounds, has predictive power for a wide range of healthcare applications, yet is currently under-explored. The main problem for those applications arises from the difficulty in collecting large labeled task-specific data for model development. Generalizable respiratory acoustic foundation models pretrained with unlabeled data would offer appealing advantages and possibly unlock this impasse. However, given the safety-critical nature of healthcare applications, it is pivotal to also ensure openness and replicability for any proposed foundation model solution. To this end, we introduce OPERA, an OPEn Respiratory Acoustic foundation model pretraining and benchmarking system, as the first approach answering this need. We curate large-scale respiratory audio datasets ($\sim$136K samples, over 400 hours), pretrain three pioneering generalizable acoustic models, and build a benchmark consisting of 19 downstream respiratory health tasks for evaluation. Our pretrained models demonstrate superior performance (against existing acoustic models pretrained with general audio on 16 out of 19 tasks) and generalizability (to unseen datasets and new respiratory audio modalities). This highlights the great promise of respiratory acoustic foundation models and encourages more studies using OPERA as an open resource to accelerate research on respiratory audio for health. The system is accessible from https://github.com/evelyn0414/OPERA.
@@ 1 Introduction
There is no open respiratory acoustic foundation model, impeding the field’s growth and understanding. Existing open-source acoustic models like AudioMAE [36] and CLAP [18] are pretrained on general audio event datasets such as YouTube audio, containing very few (around 0.3%) respiratory sounds [39, 25]. These models may not be able to effectively capture the subtle nuances of respiratory sounds, which can vary in abrupt bursts, aperiodic components, and frequency distributions, particularly across different health conditions [51]. Although a model pretrained on respiratory sounds has been recently presented [7], it is not open-source, making it hard to analyze, replicate, or compare its workings. The insights on how to effectively train generalizable respiratory acoustic models also remain limited.
@@
Figure 1: System overview of OPERA. After data curation, respiratory audio encoders are pretrained and then evaluated on various downstream health tasks.
To mitigate these gaps, in this paper, we put forward OPERA, an OPEn Respiratory Acoustic foundation model pretraining and benchmarking system (Figure 1). It curates unlabeled respiratory audio datasets, pretrain three pioneering foundational models, and evaluates them against existing pretrained acoustic models across various applications. Specifically, our contributions are:
  • We curate a unique large-scale ($\sim$136K samples, 400+ hours), multi-source (5 datasets), multi-modal (breathing, coughing, and lung sounds) and publicly available (or available on request) respiratory audio dataset for generalizable model pretraining, orders of magnitude larger than the number of respiratory audio samples in datasets used for training existing open acoustic models.
  • We pretrain 3 generalizable acoustic models with the curated unlabeled data using the most common self-supervised approaches (a contrastive learning-based transformer, a contrastive learning-based CNN model, and a generatively pretrained transformer) to study the effect of the training designs.
@@
  • We benchmark the performance of our 3 pretrained models, one commonly used acoustic feature set, and 3 open pretrained acoustic models on these tasks as a starting point for future exploration.
Extensive experiments demonstrate that our pretrained models outperform the models pretrained with general audio on 16 out of 19 benchmark tasks, confirming the power and promise of dedicated respiratory acoustic foundation models. Results also show that our models are generalizable across multiple downstream tasks, including new datasets and unseen respiratory audio modalities. This is a critical advancement towards realizing the potential of respiratory sounds as a mainstream technique for health monitoring.
@@
In summary, this paper introduces the first open-source respiratory acoustic foundation model pretraining and benchmarking system. This represents a critical first step towards comprehensive and reproducible audio foundation models for health: future foundation model research can leverage our system as an experimental resource, and application studies can take advantage of our pretrained models as feature extractors. This can facilitate progress in both machine learning and healthcare. These efforts will extend current machine learning capabilities, now able to see (via vision) and read (via natural languages), to also listen to (via audio) our health.
@@ 2.1 Pretraining in Acoustic Modeling
It is also worth noting that these open models are pretrained on general audio event datasets such as AudioSet [25], FSD50K [22], and FreeSound [23], which contains few samples of respiratory-related audio. For instance, AudioSet’s 2 million clips include only 2334 snoring, 871 cough, 834 breathing, and 1200 sneeze clips, making up only 0.3% of the total. In face of this issue, we curate large-scale respiratory audio datasets to pretrain our generalizable respiratory acoustic models for comparison.
@@
A recent work, HeAR [7], curated millions of respiratory audio clips from YouTube videos to pretrain a model using a generative SSL approach. However, neither the data nor the model are publicly available, resulting in a lack of transparency and reproducibility. Limited exploration has been conducted on the reasoning behind the chosen SSL method for various downstream tasks. Our work investigates, for the first time, open pretraining generalizable respiratory acoustic models to provide a better understanding of their limits and their potential.
@@ 3 System Overview
As shown in Figure 1, OPERA comprises three main components: data curation (including unlabeled data for pretraining and labeled data for evaluation), general-purpose pretraining to develop generalizable acoustic models (Encoder), and a benchmark comparing the pretrained models on various downstream tasks.
@@ 5.1 Benchmark Datasets and Tasks Setup
Dataset ID Task Modality #Sam. (#Sub.) Data Distribution UK COVID-19 [13] T1 Covid / Non-covid Exhalation 2500 (2500) 840 / 1660 T2 Covid / Non-covid Cough 2500 (2500) 840 / 1660 COVID-19 Sounds [71] T3 Symptomatic / Healthy Breath 4138 (3294) 2029 / 2109 T4 Symptomatic / Healthy Cough 4138 (3294) 2029 / 2109 CoughVID [49] T5 Covid / Non-covid Cough 6175 (n/a) 547 / 5628 T6 Female / Male Cough 7263 (n/a) 2468 / 4795 ICBHI [53] T7 COPD / Healthy Lung sounds 828 (90) 793 / 35 Coswara [8] T8 Smoker / Non-smoker Cough 948 (n/a) 201 / 747 T9 Female / Male Cough 2496 (n/a) 759 / 1737 KAUH [24] T10 Obstructive / Healthy Lung sounds 234 (79) 129 / 105 Respiratory@TR [3] T11 COPD severity Lung sounds 504 (42) 72 / 60 / 84 / 84 / 204 SSBPR [72] T12 Body position recognition Snoring 7468 (20) 1638 / 1454 / 1269 / 1668 / 1439 MMlung [46] T13 FVC Deep breath 40 (40) 3.402 $\pm$ 1.032 L T14 FEV1 Deep breath 40 (40) 2.657 $\pm$ 0.976 L T15 FEV1/FVC Deep breath 40 (40) 0.808 $\pm$ 0.190 L T16 FVC O Vowels 40 (40) 3.402 $\pm$ 1.032 L T17 FEV1 O Vowels 40 (40) 2.657 $\pm$ 0.976 L T18 FEV1/FVC O Vowels 40 (40) 0.808 $\pm$ 0.190 L NoseMic [10] T19 Respiratory rate Breath 1297 (16) 13.915 $\pm$ 3.386 bpm
@@
Baselines. In addition to our pretrained models, we also include a commonly used acoustic feature set and three open pretrained acoustic models in this benchmark. They are Opensmile  [19] (Emobase acoutic feature set), VGGish [31] (supervised pretrained), AudioMAE [36] (self-supervised pretrained) and CLAP [18] (language-supervised pretrained). We consider these four methods as baselines to be distinguished from our pretrained models. We also pretrain these architectures with our OPERA data and results can be found in Section A.4.
@@
Table 3: Mean reciprocal ranks on task groups (higher is better). The best model within each group is highlighted in pink and the second-best is highlighted in blue (p values reported in Section A.4).
@@ 5.2 Experimental Results
RQ1. Can pretraining a foundational model with diverse unlabeled respiratory audio data lead to better performance than baselines designed for general audio?
@@
ID Task Abbr. Opensmile VGGish AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT T1 Covid (Exhale) 0.550 $\pm$ 0.015 0.580 $\pm$ 0.001 0.549 $\pm$ 0.001 0.565 $\pm$ 0.001 0.586 $\pm$ 0.008 0.551 $\pm$ 0.010 0.605 $\pm$ 0.001 ✓* T2 Covid (Cough) 0.649 $\pm$ 0.006 0.557 $\pm$ 0.005 0.616 $\pm$ 0.001 0.648 $\pm$ 0.003 0.701 $\pm$ 0.002 0.629 $\pm$ 0.006 0.677 $\pm$ 0.001 ✓* T3 Symptom (Breath) 0.571 $\pm$ 0.006 0.571 $\pm$ 0.003 0.583 $\pm$ 0.003 0.611 $\pm$ 0.006 0.603 $\pm$ 0.005 0.610 $\pm$ 0.004 0.613 $\pm$ 0.002 ✓* T4 Symptom (Cough) 0.633 $\pm$ 0.012 0.605 $\pm$ 0.004 0.659 $\pm$ 0.001 0.669 $\pm$ 0.002 0.680 $\pm$ 0.006 0.665 $\pm$ 0.001 0.673 $\pm$ 0.001 ✓* T5 Covid (Cough) 0.537 $\pm$ 0.011 0.538 $\pm$ 0.028 0.554 $\pm$ 0.004 0.599 $\pm$ 0.007 0.578 $\pm$ 0.001 0.566 $\pm$ 0.008 0.552 $\pm$ 0.003 T6 Gender (Cough) 0.677 $\pm$ 0.005 0.600 $\pm$ 0.001 0.628 $\pm$ 0.001 0.665 $\pm$ 0.001 0.795 $\pm$ 0.001 0.721 $\pm$ 0.001 0.735 $\pm$ $0$* T7 COPD (Lung) 0.579 $\pm$ 0.043 0.605 $\pm$ 0.077 0.886 $\pm$ 0.017 0.933 $\pm$ 0.005 0.855 $\pm$ 0.012 0.872 $\pm$ 0.011 0.741 $\pm$ 0.011 T8 Smoker (Cough) 0.534 $\pm$ 0.060 0.507 $\pm$ 0.027 0.549 $\pm$ 0.022 0.680 $\pm$ 0.009 0.685 $\pm$ 0.012 0.674 $\pm$ 0.013 0.650 $\pm$ 0.005 ✓* T9 Gender (Cough) 0.753 $\pm$ 0.008 0.606 $\pm$ 0.003 0.724 $\pm$ 0.001 0.742 $\pm$ 0.001 0.874 $\pm$ 0.000 0.801 $\pm$ 0.002 0.825 $\pm$ 0.001 ✓* T10 Obstructive (Lung) 0.636 $\pm$ 0.082 0.605 $\pm$ 0.036 0.616 $\pm$ 0.041 0.697 $\pm$ 0.004 0.722 $\pm$ 0.016 0.741 $\pm$ 0.014 0.703 $\pm$ 0.016 ✓* T11 COPD severity (Lung) 0.494 $\pm$ 0.054 0.590 $\pm$ 0.034 0.510 $\pm$ 0.021 0.636 $\pm$ 0.045 0.625 $\pm$ 0.038 0.683 $\pm$ 0.007 0.606 $\pm$ 0.015 ✓* T12 Position (Snoring) 0.772 $\pm$ 0.005 0.657 $\pm$ 0.002 0.649 $\pm$ 0.001 0.702 $\pm$ 0.001 0.781 $\pm$ 0.000 0.769 $\pm$ 0.000 0.742 $\pm$ 0.001 ✓*
@@
ID Task Abbr. Opensmile VGGish AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT T13 FVC (Breath) 0.985 $\pm$ 0.743 0.904 $\pm$ 0.568 0.900 $\pm$ 0.551 0.896 $\pm$ 0.542 0.924 $\pm$ 0.583 0.848 $\pm$ 0.607 0.892 $\pm$ 0.618 ✓* T14 FEV1 (Breath) 0.756 $\pm$ 0.721 0.839 $\pm$ 0.563 0.821 $\pm$ 0.590 0.840 $\pm$ 0.547 0.837 $\pm$ 0.563 0.834 $\pm$ 0.581 0.825 $\pm$ 0.560 T15 FEV1/FVC (Breath) 0.141 $\pm$ 0.185 0.131 $\pm$ 0.146 0.129 $\pm$ 0.146 0.134 $\pm$ 0.146 0.128 $\pm$ 0.140 0.132 $\pm$ 0.141 0.128 $\pm$ 0.141 ✓* T16 FVC (Vowel) 0.850 $\pm$ 0.592 0.895 $\pm$ 0.559 0.833 $\pm$ 0.588 0.883 $\pm$ 0.560 0.885 $\pm$ 0.553 0.761 $\pm$ 0.544 0.878 $\pm$ 0.550 ✓* T17 FEV1 (Vowel) 0.730 $\pm$ 0.497 0.842 $\pm$ 0.559 0.876 $\pm$ 0.561 0.859 $\pm$ 0.541 0.780 $\pm$ 0.542 0.830 $\pm$ 0.561 0.774 $\pm$ 0.554    * T18 FEV1/FVC (Vowel) 0.138 $\pm$ 0.166 0.130 $\pm$ 0.145 0.131 $\pm$ 0.141 0.137 $\pm$ 0.147 0.132 $\pm$ 0.140 0.136 $\pm$ 0.150 0.130 $\pm$ 0.138 ✓* T19 Breathing Rate 2.714 $\pm$ 0.902 2.605 $\pm$ 0.759 2.641 $\pm$ 0.813 2.650 $\pm$ 0.947 2.636 $\pm$ 0.858 2.525 $\pm$ 0.782 2.416 $\pm$ 0.885 ✓*
RQ2. Are the pretrained respiratory acoustic models generalizable to new data?
It is crucial that foundation models can generalize to new and unseen data once developed. In our benchmark, we have 12 tasks formulated from unseen datasets (Task 8-19) and unseen respiratory audio modalities (Task 12, 16-18) not used for pretraining. Notably, our respiratory acoustic foundation models demonstrate good generalization capabilities, achieving the best performance on 5 out of 5 classification tasks and 5 out of 7 regression tasks. They are able to outperform the acoustic feature set and general audio pretrained models which are supposed to exhibit generalizability. Specifically, in Table 4, Task 8-12 all have an AUROC higher than 0.68. Comparing Task 6 and Task 9 with the same prediction target, the performance on unseen data (Task 9) is comparable. Therefore, our foundation models are generalizable, likely due to the minimal assumptions made during SSL pretraining. We have additional experiments on cross-domain zero-shot performance in Section A.4.
RQ3. How to design SSL methods and model architectures of the pretrained respiratory acoustic encoders with different applications in mind?
@@
To gain deeper insight, we further use saliency maps to explicitly inspect what our models focus on in the spectrograms of unseen audio data. Figure 3 presents examples for three tasks. We observe that the OPERA-CT model tends to focus on a few local areas of the spectrogram, showing distinct saliency peaks, whereas the GT model analyzes the global distribution with a more even saliency map. Comparing Figures 3 B.1 and B.2, GT demonstrates greater ability in detecting the continuous decline in energy after breathing. This explains why CT underperforms compared to the GT variant in lung function estimation, where global patterns are more critical.
(a) Saliency maps for T2 COVID detection with a spectrogram of three coughs. OPERA-CT highlights strong gradients for low and high-frequency bins in the middle while GT spreads attention across various time frames and frequency bins, less focused on the coughs.
(b) Saliency maps for T13 FVC estimation with a spectrogram of one respiratory cycle. OPERA-CT highlights high-frequency bins in the upper-right corner, while GT focuses on decayed energy in high-frequency bins, which is more useful for this task.
(c) Saliency maps for T13 FVC estimation with a spectrogram of four respiratory cycles. Both models highlight the time frames with silence which are useful for this task.
Figure 3: Saliency maps generated by OPERA-CT and OPERA-GT on three example tasks (T2, T13, and T19). The yellow color indicates the largest gradient on the spectrogram.